If the 100kg crate of bananas hanging from the rope were to move downwards with a constant acceleration of 1.2 m/s^2. Find the tension in the rope
1
Expert's answer
2019-09-27T09:33:49-0400

The Newton's second law gives


"ma=-T+mg"

So, the tension in the rope


"T=m(g-a)=100\:\rm kg\times (9.8\: m/s^2-1.2\: m/s^2)=860\: N"